Sheridan Inc. has issued three types of debt on January 1, 2020, the start of the company’s fiscal year.

(a) $10 million, 11-year, 14% unsecured bonds, interest payable quarterly. Bonds were priced to yield 10%.
(b) $27 million par of 11-year, zero-coupon bonds at a price to yield 10% per year.
(c) $20 million, 11-year, 9% mortgage bonds, interest payable annually to yield 10%.


Prepare a schedule that identifies the following items for each bond: (1) maturity value, (2) number of interest periods over life of bond, (3) stated rate per each interest period, (4) effective-interest rate per each interest period, (5) payment amount per period, and (6) present value of bonds at date of issue. (Round stated and effective rate per period to 2 decimal places, e.g. 10.25%. Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answer to 0 decimal places e.g. 58,971.)

Unsecured
Bonds

Zero-Coupon
Bonds

Mortgage
Bonds

(1) Maturity value $ $ $
(2) Number of interest periods
(3) Stated rate per period % % %
(4) Effective rate per period % % %
(5) Payment amount per period $ $ $
(6) Present value $ $ $
Answer #1

Unsecured
Bonds

Zero-Coupon
Bonds

Mortgage
Bonds

(1)

Maturity value

$10000000

$27000000

$20000000

(2)

Number of interest periods

44

11

11

(3)

Stated rate per period

3.5

%

0

%

9

%

(4)

Effective rate per period

2.5

%

10

%

10

%

(5)

Payment amount per period

$350000

$0

$1800000

(6)

Present value

$12650348

$9463230

$18700908

Unsecured Bonds:-

Stated rate per period = 14%/4 = 3.5%

Effective rate per period = 10%/4 = 2.5%

Payment amount per period = $10000000 * 14% * 3/12 = $350000

Present value = ($350000 * PVAF 2.5%,44) + ($10000000 * PVIF 2.5%,44)

                = ($350000 * 26.50385) + ($10000000 * 0.33740) = $12650348

Zero Coupon Bond :-

Present value = ($27000000 * PVIF 10%,11)

                 = ($27000000 * 0.35049) = $9463230

Mortgage Bonds:-

Payment amount per period = $20000000 * 9% = $1800000

Present value = ($1800000 * PVAF 10%,11) + ($20000000 * PVIF 10%,11)

              = ($1800000 * 6.49506) + ($20000000 * 0.35049) = $18700908